OpenStudy (anonymous):

need help on this word equation if 5 is subtracted by 3 times the letter m, the result is -4

OpenStudy (anonymous):

5-3m=-4 -3m=-9 m=3

OpenStudy (anonymous):

thats what i though, thnks

OpenStudy (anonymous):

so i should subtract 5 from both sides but after that shouldn't it look like this 3m=-9 not -3m=-9?

OpenStudy (anonymous):

no... because -3m remains there althought 5 is subtracted

OpenStudy (anonymous):

can you please write the whole equation step by step. please. i know the answer but confused on the steps

OpenStudy (anonymous):

its just like what i write

OpenStudy (anonymous):

ok thank you